OpenStudy (katielong):

what is the result of a positive square route x a negative square route? thanks

OpenStudy (katielong):

eg. \[-\sqrt{3}\times \sqrt{3}\]

OpenStudy (apoorvk):

so, that would be the negative of the no. whose square root is being taken!

OpenStudy (anonymous):

That would be the same as \[-\sqrt{3*3}\]

OpenStudy (katielong):

arite! so that would be -3?

OpenStudy (anonymous):

That is correct!

OpenStudy (katielong):

cheers dude!! :D

OpenStudy (anonymous):

No problem :-)

OpenStudy (unklerhaukus):

\[\sqrt{3}\times\sqrt3=\sqrt{3\times3}=3\]\[-\sqrt{3}\times\sqrt3=-\sqrt{3\times3}=-3\]\[-\sqrt{3}\times-\sqrt3=\sqrt{3\times3}=3\] \[\sqrt{-3}\times\sqrt3=\sqrt{-1}\times\sqrt{3\times3}=3i\]\[\sqrt{-3}\times\sqrt{-3}=\sqrt{-3\times-3}=\sqrt{3\times 3}=3\]